# Pump-efficient Josephson parametric amplifiers with high saturation power

### v1

@@ Abstract
Circuit QED based quantum information processing relies on low noise amplification for signal readout. In the realm of microwave superconducting circuits, this amplification is often achieved via Josephson parametric amplifiers (JPA). In the past, these amplifiers exhibited low power added efficiency (PAE), which is roughly the fraction of pump power that is converted to output signal power. This is increasingly relevant because recent attempts to build high saturation power amplifiers achieve this at the cost of very low PAE, which in turn puts a high heat load on the cryostat and limits the number of these devices that a dilution refrigerator can host. Here, we numerically investigate upper bounds on PAE. We focus on a class of parametric amplifiers that consists of a capacitor shunted by a nonlinear inductive block. We first set a benchmark for this class of amplifiers by considering nonlinear blocks described by an arbitrary polynomial current-phase relation. Next, we propose two circuit implementations of the nonlinear block. Finally, we investigate chaining polynomial amplifiers. We find that while amplifiers with higher gain have a lower PAE, regardless of the gain there is considerable room to improve as compared to state of the art devices. For example, for a phase-sensitive amplifier with a power gain of 20 dB, the PAE is $\sim 0.1\%$ for typical JPAs, 5.9% for our simpler circuit JPAs, 34% for our more complex circuit JPAs, 48% for our arbitrary polynomial amplifiers, and at least 95% for our chained amplifiers.
@@ I Introduction
The Josephson parametric amplifier (JPA) plays an important role in superconducting quantum computing. It uses the Josephson junction to amplify weak microwave signals with nearly quantum-limited added noise. Due to its ultralow-noise performance, the JPA is vital for readout [1, 2, 3, 4] for applications such as quantum state tomography [5], and, crucially, quantum error correction [6, 7, 8, 9].
A critical characteristic of the JPA is its ability to add the minimum noise throughout the parametric amplification process with a power gain of around 20 dB [10, 11, 12]. These noise and gain characteristics make JPAs suitable as the first stage of an amplifier chain, with the gain requirement being the minimum value to strongly saturate commercial cryogenic high electron mobility transistor (HEMT) amplifiers [12, 13]. Parametric amplification happens when the circuit’s parameter (e.g., inductance) is varied periodically at a specific frequency. In a microwave-pumped JPA, the periodic variation is achieved by using the nonlinearity of the Josephson junction to couple in a strong pump wave. Specifically, the nonlinearity mixes different frequency components; the pump wave at frequency $\omega_{\text{p}}$ and the signal wave at frequency $\omega_{\text{s}}$ mix and generate an idler wave at frequency $\omega_{\text{i}}=\omega_{\text{p}}-\omega_{\text{s}}$. The parametric process transfers energy from the pump wave to the signal and idler waves and thus amplifies the incoming signals [14, 15, 16].
Based on signal and idler frequencies, we usually classify the parametric amplifiers into phase-sensitive and phase-preserving amplifiers [12]. The phase-sensitive amplifier is also called the degenerate amplifier because its signal and idler frequencies are hosted in the same physical mode. Since the pump frequency is exactly twice the mode center frequency in this case, the power gain of the degenerate amplifier is sensitive to the relative phase between the pump and signal waves near the center of the amplifier’s mode frequency. In contrast, the phase-preserving amplifier, also known as the nondegenerate amplifier, has different signal and idler frequencies. In this case, the amplitude and phase information of the signal wave is maintained during the amplification process.
@@
The subject of this paper is pump efficiency, which characterizes the capacity of an amplifier to convert input power from the pump to output signal power without distorting the signal. Specifically, we use the power added efficiency (PAE) [23, p. 597], which is defined as
@@
where $G(A_{\text{s}})$ is the power gain as a function of the input signal amplitude $A_{\text{s}}$, $A_{\text{p}}$ is the amplitude of the pump, and $\omega_{\text{s}}$ and $\omega_{\text{p}}$ refer to the signal and pump frequencies, respectively. Here, $P_{\text{in}}$ is the power of the input signal and $P_{\text{pump}}$ is the power of the pump wave. We note that $P_{\text{out}}$ consists only of the power being converted as output at the signal frequency, while ignoring the power at the idler frequency. Therefore, for phase-preserving amplifiers, but not phase-sensitive amplifiers, approximately half of the output power, which comes out at the idler frequency, is not accounted for in the PAE. We choose this definition because we expect that only the signal frequency will be used downstream. We also note that in this work, we will focus on amplifiers with a target gain of $G_{\text{t}}=20~{}\text{dB}$. Finally, we note that for convenience of the reader, we list commonly used symbols in Tab. 3.
@@
In this paper, we will investigate the ultimate limit on the PAE of a class of parametric amplifiers illustrated in Fig. 1. We will show that amplifiers within this class can have a PAE orders of magnitude higher than typical JPA designs. Specifically, we consider amplifier circuits of the general form shown in Fig. 1 that consist of a capacitor (denoted by $C$) in parallel with a nonlinear inductive block with energy $E[\phi]$. We first use this framework to explore the ultimate limit on PAE by considering arbitrary functions $E[\phi]$. Next, we explore buildable circuits where $E[\phi]$ represents the energy of a collection of inductors and Josephson junctions. Finally, we explore improving PAE by using amplifier chains. In the next three paragraphs, we summarize the results of these three directions.
@@
where $L_{\text{eff}}$ is the effective inductance of the inductive block. We tune the $g_{i}$ coefficients in this polynomial amplifier in order to maximize saturation power and, consequently, PAE. Searching through the space of polynomial amplifiers sets a bound on what level of efficiency is achievable given a certain order of nonlinearity, i.e., we can increase the PAE by adding further terms to $E[\phi]$. We have explored up to ninth order amplifiers. In Sec. III.1, we find that the PAE tends to saturate at around seventh order, where the maximum achievable PAE is at least 48% for degenerate amplifiers with $G_{\text{t}}=20~{}\text{dB}$.
Next, we investigate whether it is possible to achieve the same PAE from buildable amplifiers with inductive blocks composed of inductors and Josephson junctions. We replace the generic $E[\phi]$ by the energy of a circuit composed of these circuit elements. We choose circuits which have a comparable number of tuning parameters as in the polynomial case and adjust these parameters to optimize saturation power. In Sec. III.2, we show that we can reach a maximum PAE of 34% for degenerate amplifiers with $G_{\text{t}}=20~{}\text{dB}$.
Finally, we explore increasing PAE by chaining amplifiers. The PAE of an amplifier usually decreases as the gain increases. Therefore, one might expect to obtain an amplifier with a higher PAE by connecting several smaller gain amplifiers together. In Sec. III.3, we find that we can increase the PAE at $20$ dB from around $50\%$ to more than $95\%$ by chaining $\sim 30$ amplifiers (see in Fig. 21).
@@ II Amplifier Model and Numerics
In this section, we lay out the model for a class of parametric amplifiers composed of a capacitor shunted by a nonlinear inductive block (shown in Fig. 2), derive the classical input-output relation, and obtain the equation of motion (EOM). Our calculations mainly follow Ref. 24.
@@ II.2 Equations of motion and input-output relations
where ${}^{\prime}$ denotes the spatial derivative. We define the incoming and outgoing waves $\phi_{\text{in}}$ and $\phi_{\text{out}}$ such that
@@ II.2.1 Polynomial amplifiers
A polynomial amplifier of order $n$ is defined by a polynomial energy $E[\phi]=\frac{1}{2L_{\text{eff}}}\phi^{2}+g_{3}\phi^{3}+\cdots+g_{n}\phi^{n}$. The corresponding current-phase relation $J[\phi]$ is a polynomial of order $n-1$,
@@
where $\omega^{2}_{0}=\frac{1}{CL_{\text{eff}}}$ is the natural frequency of the amplifier. We note that in this polynomial amplifier, the third order ($g_{3}$) term is primarily responsible for amplification. However, the third order term dynamically generates higher order nonlinear terms which limit the amplifier’s performance at high signal power [25, 21]. We aim to cancel the dynamically generated terms with higher order static terms in Eq. 14 in order to increase the saturation power and hence the PAE of the amplifier at fixed pump amplitude.
@@ II.2.2 Josephson parametric amplifiers
We represent the inductive block $E[\varphi]$ by a network of internal inductive components which make up our circuit design. Specifically, the circuit is composed of a set of internal nodes $\varphi_{i}$ that are connected by the inductive components. The inductive block does not contain any explicit dynamic terms since parasitic capacitances in this block are small and can be ignored. They are, however, sufficiently large that we operate in the singular regime of the Lagrangian, and thus we can neglect quantum effects as discussed in Ref. 26.
For each node $\varphi_{i}$, we write that $\frac{\partial\mathcal{L}}{\partial\varphi_{i}}=J_{i}=J_{\text{ext},i}$ where $J_{\text{ext},i}$ is the current bias at node $\varphi_{i}$. Typically, nodes are not current biased and thus have $J_{i}=0$. From here we proceed as in Ref. 21.
We also bias the circuit by applying external magnetic flux through closed loops in the circuit. Without loss of generality, we model magnetic flux bias by describing it as a phase offset across one of the Josephson junctions in the loop. In particular, a Josephson junction in a loop with an applied external flux of $\Phi_{\text{ext}}$ will have a contribution of $\frac{\phi_{0}^{2}}{2L}(\varphi_{1}-\varphi_{2}+\Phi_{\text{ext}})^{2}$ to the Lagrangian.
@@
$\ddot{\varphi}+K\dot{\varphi}+\frac{\varphi}{nLC}+\frac{i_{\text{c}}}{C}\sin\left(\frac{\varphi}{n}+\Phi_{\text{ext}}\right)=2K\dot{\varphi_{\text{in}}},\end{aligned}$ (17)
where $L$, $i_{\text{c}}$, and $\Phi_{\text{ext}}$ are the linear inductance, Josephson critical current, and external flux through each RF-SQUID, respectively. In general, we can construct the current-phase relation $J[\varphi]$ for the designed amplifier circuit and apply it to the EOM specified in equation (12).
Figure 3: Amplifier circuit with an inductive block composed of a chain of $n$-many RF-SQUIDs. This type of circuit can have high saturation power, but it is limited to low PAE.
@@ II.3 Numerical solutions of EOM and optimization algorithm for PAE
where $\omega_{\text{s}},\omega_{\text{p}}$ are the signal and pump frequencies, $A_{\text{s}},A_{\text{p}}$ are the signal and pump amplitudes, and $\delta$ is the phase difference between the signal and pump wave. Because the differential equation (14) is nonlinear, inhomogeneous and non-autonomous, obtaining analytical solutions is not possible in general. Thus, we first introduce two numerical methods, the direct time integration (DTI) and the harmonic balance (HB) method, that we use to solve this differential equation.
The DTI method is based on numerical integration techniques that approximate the differential equation by using finite differences. The DTI method uses function values and finite differences at current and previous time intervals to find the value at the next interval. In our calculation, we use the Python method scipy.integrate.odeint to solve the nonlinear differential equation [29]. Since we want to obtain a steady-state solution, we choose our integration time to be at least 2000 periods of the signal wave. Then, to remove the initial transient, we only analyze the data from the last quarter of the integration time. After obtaining the solution in the time domain, we apply the Discrete Fourier Transform (DFT) to isolate the outgoing signal wave from the rest of the solution. This requires the last quarter of the integration time to be a multiple of $T_{\min}=\gcd(\frac{2\pi}{\omega_{\text{s}}},\frac{2\pi}{\omega_{\text{p}}})$ which dictates our choice of $\omega_{\text{p}}/\omega_{\text{s}}$. Finally, the PAE is determined by calculating both the saturation power gain and the amplitude of the signal wave at that gain.
@@
$\phi(t)=a_{0}+a_{1}e^{i\omega_{0}t}+\cdots+a_{n}e^{in\omega_{0}t}+h.c..$ (19)
Then, the solution can be obtained by matching coefficients on both sides of Eq. (14). To ensure we have chosen a sufficiently large $n$, we also compare the HB method with the DTI method. For degenerate amplifiers, the HB method is typically much faster than the DTI method. On the other hand, for nondegenerate amplifiers, we need to consider many more harmonics in the HB method. In this case, the DTI method is faster, and is the one we use. A more comprehensive analysis of these methods is provided in Appendix B.
@@
with the incoming wave $\phi_{\text{in}}=A_{\text{s}}\sin(\omega_{\text{s}}t)+A_{\text{p}}\sin(\omega_{\text{p}}t+\delta)$. First, we rescale time $t\to t/\omega_{0}$, i.e. we work in units where $\omega_{0}=1$. Next, we scale flux so that $A_{\text{p}}$ is normalized to be $0.5$. In this case, the pump power is $\omega_{\text{p}}^{2}$, and the PAE is $4(G-1)A_{\text{s}}^{2}\omega_{\text{s}}^{2}/\omega_{\text{p}}^{2}$. Finally, we rescale coefficients $g_{i}$ to absorb $1/C$. In the following, we will use the same letter to denote the normalized polynomial coefficients. After rescaling, our differential equation becomes
$\ddot{\phi}+K{\dot{\phi}}+{\phi}+3g_{3}\phi^{2}+\cdots+ng_{n}\phi^{n-1}=2K\dot{\phi}_{\text{in}},$ (21)
with the incoming wave $\phi_{\text{in}}=A_{\text{s}}\sin(\frac{\omega_{\text{s}}}{\omega_{0}}t)+\frac{1}{2}\sin(\frac{\omega_{\text{p}}}{\omega_{0}}t+\delta)$.
@@
While we wish to maximize the PAE of the amplifier directly, this is difficult, and therefore we use a proxy. Specifically, we minimize the difference between the amplifier gain and $G_{\text{t}}$ over the largest possible range of $A_{\text{s}}$. We take the highest PAE below the saturation power of the amplifier as the amplifier’s PAE. This maximum PAE typically occurs at or near the saturation power of the amplifier (see in Fig. 4). The optimization variables we use are $K,g_{3},g_{4},\cdots$. In the remainder of this section, we will present the algorithm step by step with necessary explanations. A pictorial demonstration of the algorithm is also provided in Fig. 5.
@@
Figure 4: Power gain and PAE versus dimensionless signal amplitude $A_{\text{s}}$ of a 7-th order 20 dB degenerate amplifier. The parameters of this amplifier can be found in Tab. 4. The saturation amplitude $A_{\text{sat}}$ is the signal amplitude where the gain curve intersects the $G_{\text{t}}\pm 1$ dB lines. The PAE of an amplifier is defined to be the maximum PAE below the saturation power, which typically appears at or around $A_{\text{sat}}$.
@@
Step 1. Choose an initial range $[0,R]$ for $A_{\text{s}}$, and a set of initial variables $\{K^{(0)}$, $g_{3}^{(0)}$, $g_{4}^{(0)},\cdots\}$. This set of variables generates an initial gain curve as shown in Fig. 5(1).
Step 2. Divide the region $[0,R]$ into $N$ sub-intervals and define the cost function $f$ as the sum of the squared distance as illustrated in Fig. 5(2).
@@
$&f\left(K^{(0)},g_{3}^{0},g_{4}^{(0)},\cdots\right)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\\ &\sum_{m=1}^{N}\left(G\left(A_{\text{s}}=\frac{mR}{N},K^{(0)},g_{3}^{(0)},g_{4}^{(0)},\cdots\right)-G_{\text{t}}\right)^{2}.\end{aligned}$ (22)
Step 3. Minimize the cost function $f$ using the gradient descent algorithm. By doing so, we will get a new collection of coefficients $\{K^{(1)}$, $g_{3}^{(1)}$, $g_{4}^{(1)},\cdots\}$.
@@
Step 4. Calculate the saturation point for the new curve, which is shown in Fig. 5(4), and reset the new range to be $[0,A_{\text{sat}}]$ if $A_{\text{sat}}$ exceeds the previous range $R$; otherwise, stop the optimization.
@@
Step 5. Redefine the cost function $f$ within the new range as depicted in Fig. 5(5), and repeat steps 3 to 5 until the optimization stops.
@@
Step 6. Record the optimized values of the amplifier parameters $K,g_{3},g_{4},\cdots$.
@@
The solution and optimization of amplifiers described by circuits is similar to the description above, and we provide details of this procedure in Appendix C.

### v2

@@ Abstract
Circuit QED based quantum information processing relies on low noise amplification for signal readout. In the realm of microwave superconducting circuits, this amplification is often achieved via Josephson parametric amplifiers (JPA). In the past, these amplifiers exhibited low power added efficiency (PAE), which is roughly the fraction of pump power that is converted to output signal power. This is increasingly relevant because recent attempts to build high saturation power amplifiers achieve this at the cost of very low PAE, which in turn puts a high heat load on the cryostat and limits the number of these devices that a dilution refrigerator can host. Here, we numerically investigate upper bounds on PAE. We focus on a class of parametric amplifiers that consists of a capacitor shunted by a nonlinear inductive block. We first set a benchmark for this class of amplifiers by considering nonlinear blocks described by an arbitrary polynomial current-phase relation. Next, we identify that it is important for amplifiers with inductive blocks composed of repeating elements to have monotonic current-phase relations for each element in order to avoid exciting high-frequency modes. Using this design rule, we propose two circuit implementations for repeating elements in JPA inductive blocks. Finally, we investigate polynomial amplifier chains. We find that while amplifiers with higher gain have a lower PAE, regardless of the gain there is considerable room to improve as compared to state of the art devices. For example, for a degenerate amplifier with a power gain of 20 dB, the PAE is $\sim 0.1\%$ for typical JPAs, 37.9% for our simpler circuit JPAs, 42.6% for our more complex circuit JPAs, 63.3% for our arbitrary polynomial amplifiers, and at least 98% for our amplifier chains.
{CJK*}
UTF8bsmi
@@ I Introduction
The Josephson parametric amplifier (JPA), which has been extensively studied for its low noise profile [1, 2, 3, 4], plays an important role in superconducting quantum computing. It uses the Josephson junction to amplify weak microwave signals with nearly quantum-limited added noise. Due to its ultralow-noise performance, the JPA is vital for qubit state readout [5, 6, 7, 8] for applications such as quantum state tomography [9], and, crucially, quantum error correction [10, 11, 12, 13].
A critical characteristic of the JPA is its ability to add the minimum noise throughout the parametric amplification process with a power gain of around 20 dB [14, 15, 16]. These noise and gain characteristics make JPAs suitable as the first stage of an amplifier chain, with the gain requirement being the minimum value to saturate the excess noise of commercial cryogenic high electron mobility transistor (HEMT) amplifiers [16, 17]. Parametric amplification happens when the circuit’s parameter (e.g., inductance) is varied periodically at a specific frequency. In a microwave-pumped JPA, the periodic variation is achieved by using the nonlinearity of the Josephson junction to couple in a strong pump wave. Specifically, the nonlinearity mixes different frequency components of microwave signals. In particular, we consider the three-wave mixing case, where the pump wave at frequency $\omega_{\text{p}}$ and the signal wave at frequency $\omega_{\text{s}}$ mix and generate an idler wave at frequency $\omega_{\text{i}}=\omega_{\text{p}}-\omega_{\text{s}}$. The parametric process transfers energy from the pump wave to the signal and idler waves and thus amplifies the incoming signals [18, 19, 20].
Based on signal and idler frequencies, parametric amplifiers are classified into phase-sensitive and phase-preserving amplifiers [16]. The phase-sensitive amplifier is also called the degenerate amplifier because its signal and idler frequencies are hosted in the same physical mode. Since the pump frequency is exactly twice the mode center frequency in this case, the power gain of the degenerate amplifier is sensitive to the relative phase between the pump and signal waves near the center of the amplifier’s mode frequency. In contrast, the phase-preserving amplifier, also known as the nondegenerate amplifier, has different signal and idler frequencies. In this case, the amplitude and phase information of the signal wave is maintained during the amplification process.
@@
The subject of this paper is pump efficiency, which characterizes the capacity of an amplifier to convert input power from the pump to output signal power without distorting the signal. Specifically, we use the power added efficiency (PAE) as given on page 597 in Ref. [27], which is defined as
@@
where $G(A_{\text{s}})$ is the power gain as a function of the input signal amplitude $A_{\text{s}}$, $A_{\text{p}}$ is the amplitude of the pump, and $\omega_{\text{s}}$ and $\omega_{\text{p}}$ refer to the signal and pump frequencies, respectively. Here, $P_{\text{in}}$ is the power of the input signal and $P_{\text{pump}}$ is the power of the pump wave. We note that $P_{\text{out}}$ consists only of the power being converted as output at the signal frequency, while ignoring the power at the idler frequency. Therefore, for nondegenerate amplifiers, but not degenerate amplifiers, approximately half of the output power, which comes out at the idler frequency, is not accounted for in the PAE. We choose this definition because we expect that only the signal frequency will be used downstream. We also note that in this work, we will focus on amplifiers with a target gain of $G_{\text{t}}=20~{}\text{dB}$. Finally, we note that for convenience of the reader, we list commonly used symbols in Tab. 3.
@@
In this paper, we will investigate the ultimate limit on the PAE of a class of parametric amplifiers illustrated in Fig. 1. We will show that amplifiers within this class can have a PAE orders of magnitude higher than typical JPA designs. Specifically, we consider amplifier circuits of the general form shown in Fig. 1 that consist of a capacitor (denoted by $C$) in parallel with a nonlinear inductive block with energy $E[\phi]$  [^1] Note that we do not consider a coupling capacitor to the transmission line, as is commonly used in experiment. This is because coupling capacitors reduce effective PAE and we are interested in the PAE of the amplifier after the plane of the coupling capacitor.. We first use this framework to explore the ultimate limit on PAE by considering arbitrary functions $E[\phi]$. Next, we explore buildable circuits where $E[\phi]$ represents the energy of a collection of inductors and Josephson junctions. We also introduce a rule for designing amplifier circuits – namely that the current-phase relation of the inductive block should be monotonic. Finally, we explore improving PAE by using amplifier chains. In the next three paragraphs, we summarize the results of these three directions.
@@
where $L_{\text{eff}}$ is the effective inductance of the inductive block. We tune the $g_{i}$ coefficients in this polynomial amplifier in order to maximize saturation power and, consequently, PAE. Searching through the space of polynomial amplifiers sets a bound on what level of efficiency is achievable given a certain order of nonlinearity, i.e., we can increase the PAE by adding further terms to $E[\phi]$. We have explored up to tenth order amplifiers. In Subsec. III.1, we find that the PAE tends to saturate at around sixth order, where the maximum achievable PAE is at least 63% for degenerate amplifiers with $G_{\text{t}}=20~{}\text{dB}$.
Next, we investigate whether it is possible to achieve the same PAE from buildable amplifiers with inductive blocks composed of inductors and Josephson junctions. We replace the generic $E[\phi]$ by the energy of a circuit composed of these circuit elements. We choose circuits which have a comparable number of tuning parameters as in the polynomial case and adjust these parameters to optimize saturation power. In Subsec. III.3, we show that we can reach a maximum PAE of 42.6% for degenerate amplifiers with $G_{\text{t}}=20~{}\text{dB}$.
Finally, we explore increasing PAE by chaining together amplifiers. The PAE of an amplifier usually decreases as the gain increases. Therefore, one might expect to obtain an amplifier with a higher PAE by connecting several smaller gain amplifiers together. In Subsec. III.4, we find that we can increase the PAE at $20$ dB from around $60\%$ to more than $98\%$ by chaining $\sim 15$ amplifiers (see in Fig. 21).
@@ II Amplifier Model and Numerics
In this section, we lay out the model for a class of parametric amplifiers composed of a capacitor shunted by a nonlinear inductive block (shown in Fig. 2), derive the classical input-output relation, and obtain the equation of motion (EOM). Our calculations mainly follow Ref. 28.
@@ II.2 Equations of motion and input-output relations
where denotes the spatial derivative. We define the incoming and outgoing waves $\phi_{\text{in}}$ and $\phi_{\text{out}}$ such that
@@ II.2.1 Polynomial amplifiers
A polynomial amplifier of order $n$ is defined by a polynomial energy,
$E[\phi]=\frac{1}{2L_{\text{eff}}}\phi^{2}+g_{3}\phi^{3}+\cdots+g_{n}\phi^{n}.$ (13)
To ensure stability of the amplifier, we require $n=2m$ to be a even number and $g_{n}>0$, to guarantee the existence of a global energy minimum. Ideally, we want $\phi=0$ to be the unique minimum point of the energy function, but this is often impractical. Instead, we demand that in the vicinity of $\phi=0$, $E[\phi]$ is a monotonically increasing function of $\phi$ as we move away from the origin in either direction.
The corresponding current-phase relation $J[\phi]$ is a polynomial of order $n-1$,
@@
where $\omega^{2}_{0}=\frac{1}{CL_{\text{eff}}}$ is the natural frequency of the amplifier. We note that in this polynomial amplifier, the third order ($g_{3}$) term is primarily responsible for amplification. However, the third order term dynamically generates higher order nonlinear terms which limit the amplifier’s performance at high signal power [29, 25]. We aim to cancel the dynamically generated terms with higher order static terms in Eq. (15) in order to increase the saturation power and hence the PAE of the amplifier at fixed pump amplitude.
@@ II.2.2 Josephson parametric amplifiers
We represent the inductive block $E[\varphi]$ by a network of internal inductive components which make up our circuit design. Specifically, the circuit is composed of a set of internal nodes $\varphi_{i}$ that are connected by the inductive components. In Ref. 30, it was shown that stray capacitances could have important effects. We discuss the effect of capacitive terms within the inductive block in Subsec. III.2.
For each node $\varphi_{i}$, we write that $\frac{1}{\phi_{0}}\frac{\partial\mathcal{L}}{\partial\varphi_{i}}=J_{i}=J_{\text{ext},i}$ where $J_{\text{ext},i}$ is the current bias at node $\varphi_{i}$. Typically, nodes are not current biased and thus have $J_{i}=0$. From here we proceed as in Ref. 25.
We also bias the circuit by applying external magnetic flux through closed loops in the circuit. Without loss of generality, we model magnetic flux bias by describing it as a phase offset across one of the Josephson junctions in the loop. In particular, a Josephson junction in a loop with an applied external flux of $\Phi_{\text{ext}}$ will have a contribution of $\frac{\phi_{0}^{2}}{2L}(\varphi_{1}-\varphi_{2}+\varphi_{\text{ext}})^{2}$ to the Lagrangian, where $\varphi_{\text{ext}}=\Phi_{\text{ext}}/\phi_{0}$.
@@
$\ddot{\varphi}+K\dot{\varphi}+\frac{\varphi}{nLC}+\frac{i_{\text{c}}}{\phi_{0}C}\sin\left(\frac{\varphi}{n}+\varphi_{\text{ext}}\right)=2K\dot{\varphi_{\text{in}}},\end{aligned}$ (18)
where $L$, $i_{\text{c}}$, and $\varphi_{\text{ext}}$ are the linear inductance, Josephson critical current, and dimensionless external flux through each RF-SQUID, respectively. In general, we can construct the current-phase relation $J[\varphi]$ for the designed amplifier circuit and apply it to the EOM specified in Eq. (12).
Figure 3: Amplifier circuit with an inductive block composed of a chain of $n$-many RF-SQUIDs.
@@ II.3 Numerical solutions of EOM and optimization algorithm for PAE
where $\omega_{\text{s}},\omega_{\text{p}}$ are the signal and pump frequencies, $A_{\text{s}},A_{\text{p}}$ are the signal and pump amplitudes, and $\delta$ is the phase difference between the signal and pump wave. Because the differential equation (15) is nonlinear, inhomogeneous and non-autonomous, obtaining analytical solutions is not possible in general. Thus, we first introduce two numerical methods, the direct time integration (DTI) and the harmonic balance (HB) method, which we use to solve this differential equation.
The DTI method is based on numerical integration techniques that approximate the differential equation by using finite differences. The DTI method uses function values and finite differences at current and previous time intervals to find the value at the next interval. In our calculation, we use the Python method scipy.integrate.odeint [33] to solve the nonlinear differential equation. Since we want to obtain a steady-state solution, we choose our integration time to be at least 2000 periods of the signal wave. Then, to remove the initial transient, we only analyze the data from the last quarter of the integration time. After obtaining the solution in the time domain, we apply the discrete Fourier transform (DFT) to isolate the outgoing signal wave from the rest of the solution. This requires that the last quarter of the integration time is a multiple of $T_{\min}=\text{lcm}(\frac{2\pi}{\omega_{\text{s}}},\frac{2\pi}{\omega_{\text{p}}})$, where lcm is a function giving the least common multiple of its inputs, which dictates our choice of $\omega_{\text{p}}/\omega_{\text{s}}$. Finally, the PAE is determined by calculating both the saturation power gain and the amplitude of the signal wave at that gain.
@@
$\phi(t)=a_{0}+a_{1}e^{i\omega_{\text{s}}t}+\cdots+a_{n}e^{in\omega_{\text{s}}t}+h.c..$ (20)
Then, the solution can be obtained by matching coefficients on both sides of Eq. (15). To ensure that we have chosen a sufficiently large $n$, we also compare the HB method with the DTI method. For degenerate amplifiers, the HB method is typically much faster than the DTI method. On the other hand, for nondegenerate amplifiers, we need to consider many more harmonics in the HB method. In this case, the DTI method is faster, and is the one we use. A more comprehensive analysis of these methods is provided in Appendix B.
@@
with the incoming wave $\phi_{\text{in}}=A_{\text{s}}\sin(\omega_{\text{s}}t)+A_{\text{p}}\sin(\omega_{\text{p}}t+\delta)$. First, we rescale time $t\to 2t/\omega_{\text{p}}$, i.e., we work in units where $\omega_{\text{p}}=2$. Next, we scale flux so that $A_{\text{p}}$ is normalized to be $0.5$. In this case, the pump power is $A_{\text{p}}^{2}\omega_{\text{p}}^{2}=1$ and the PAE is $(G-1)\omega_{\text{s}}^{2}A_{\text{s}}^{2}$. Finally, we rescale coefficients $g_{i}$ to absorb $1/C$. In the following, we will use the same letter to denote the normalized polynomial coefficients. After rescaling, our differential equation becomes
$\ddot{\phi}+K{\dot{\phi}}+\omega_{0}^{2}{\phi}+3g_{3}\phi^{2}+\cdots+ng_{n}\phi^{n-1}=2K\dot{\phi}_{\text{in}},$ (22)
with the incoming wave $\phi_{\text{in}}=A_{\text{s}}\sin(\omega_{\text{s}}t)+\frac{1}{2}\sin(2t+\delta)$.
@@
While we wish to maximize the PAE of the amplifier directly, this is difficult, and therefore we use a proxy. Specifically, we minimize the difference between the amplifier gain and $G_{\text{t}}$ over the largest possible range of $A_{\text{s}}$. We take the highest PAE below the saturation power of the amplifier as the amplifier’s PAE. This maximum PAE typically occurs at or near the saturation power of the amplifier (see Fig. 4). The optimization variables we use are $\omega_{0},K,g_{3},g_{4},\cdots$. We use $\omega_{0}$ as an optimization variable since it is possible that the amplifier should have a natural frequency which is different than the target signal frequency $\omega_{s}$ [29]. In the remainder of this section, we will present the algorithm step by step with necessary explanations. A pictorial demonstration of the algorithm is also provided in Fig. 5.
@@
Figure 4: Power gain and PAE versus dimensionless signal amplitude $A_{\text{s}}$ of a 6-th order 20 dB degenerate amplifier. The parameters of this amplifier can be found in Tab. 4. The saturation amplitude $A_{\text{sat}}$ is the signal amplitude where the gain curve intersects the $G_{\text{t}}\pm 1$ dB lines. The PAE of an amplifier is defined to be the maximum PAE below the saturation power, which typically appears at or around $A_{\text{sat}}$.
@@
Step 1. Choose an initial range $[0,R]$ for $A_{\text{s}}$, and a set of initial variables $\{\omega_{0}^{(0)},K^{(0)}$, $g_{3}^{(0)}$, $g_{4}^{(0)},\cdots\}$. This set of variables generates an initial gain curve as shown in Fig. 5(a).
Step 2. Divide the region $[0,R]$ into $N$ sub-intervals and define the cost function $f$ as the sum of the squared distance (illustrated in Fig. 5(b)) plus the penalty that ensures that $E[\phi]$ is a monotonically increasing function of $\phi$ as one moves away from $\phi=0$ in either direction. For finer optimization, we use an alternative cost functions. For more information on this and the applied penalty, see Appendix C. We have checked that the range over which we enforce the condition on $E[\phi]$ is larger than the largest fluctuations of $\phi(t)$ that we observe.
@@
$&f\left(\omega_{0}^{(0)},K^{(0)},g_{3}^{0},\cdots\right)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\\ &\sum_{m=1}^{N}\left(G\left(A_{\text{s}}=\frac{mR}{N},\omega_{0}^{(0)},K^{(0)},g_{3}^{(0)},\cdots\right)-G_{\text{t}}\right)^{2}+\text{penalty}\end{aligned}$ (23)
[…]
@@
Step 4. Calculate the saturation point for the new curve, which is shown in Fig. 5(d), and reset the new range to be $[0,A_{\text{sat}}]$ if $A_{\text{sat}}$ exceeds the previous range $R$; otherwise, stop the optimization.
@@
Step 5. Redefine the cost function $f$ within the new range as depicted in Fig. 5(e), and repeat steps 3 to 5 until the optimization stops.
@@
Step 6. Record the optimized values of the amplifier parameters $\omega_{0},K,g_{3},\cdots$.
@@
The solution and optimization of amplifiers described by circuits is similar to the description above, and we provide details of this procedure in Appendix D.